A bee flies on a trajectory such that its polar coordinate at time t are given by "r=bt\/T"2(2T-t) "\theta=t/T" (0<t<2T) where b and T are positive constants. Find the velocity vector of the bee at time t. Show that the least speed achieved by the bee is b/T. Find the acceleration of the bee at this instant.


1
Expert's answer
2021-10-05T15:12:53-0400
"r=\dfrac{bt}{T^2}(2T-t), \theta=\dfrac{t}{T}, (0<t<2T)"

"\dot{r}=\dfrac{2b}{T^2}(T-t), \dot{\theta}=\dfrac{1}{T}"

"\ddot{r}=-\dfrac{2b}{T^2}, \ddot{\theta}=0"

"v=\dfrac{2b}{T^2}(T-t)\hat{r}+\dfrac{bt}{T^3}(2T-t)\hat{\theta}"

"a=(-\dfrac{2b}{T^2}-\dfrac{bt}{T^4}(2T-t))\hat{r}+\dfrac{4b}{T^3}(T-t)\hat{\theta}"